If Veronica Vaughn spends all of her daily income on cigarettes and Yoo Hoo, she can afford 10 packs of cigarettes and 10 bottles of Yoo Hoo. She can also afford 6 packs of cigarettes and 22 bottles of Yoo Hoo.



1. What is the relative price of cigarettes in terms of bottles of the yoo - hoo ( ie what is the ratio of prices Pc/ Py)




2. Exactly how much income does Veronica earn in one week if the price of cigarettes is $6? Write down the budget equation for Veronica that is a function of the packs of cigarettes C,and the number of bottle of yoo- hop, Y.

1
Expert's answer
2021-11-17T10:00:31-0500

a.). Budget constraint: PcC + PyY = M


M = Income


10Pc + 10Py = M


6Pc + 22Py = M


4Pc – 2Py = M


4Pc = 2Py


2Pc = Py


Py = 2Pc


Price ratio = "\frac{Pc}{Py} = \frac{Pc}{2Pc} = 2"


b.). PcC + PyY = M


Pc = 6


60 + 10Py = M


36 + 22Py = M


24 – 12Py = 0


24 = 12Py


Py = 2


M = PcC + PyY


M = 6(10) + 2(10) = 60 + 20 = 80


Income (M) = 80